# Towards the understanding of heavy quarks hadronization: from leptonic to heavy-ion collisions

### v2

@@ II.2.5 EPOS4(HQ)
Historically, the scattering modelling in EPOS was based on a zero-momentum limit of cut pomerons and hence did not aim at describing high-$p_{\perp}$ jet fragmentation (or other perturbative scattering processes). This has changed with EPOS4, which incorporates both factorization for hard processes and saturation for the low-$x$ scatterings. More specifically related to the purposes of this review, the main aspect of EPOS4HQ is the inclusion of the scattering dynamics of heavy quarks in the “corona” medium even in pp collisions and of hadronization via coalescence plus fragmentation [101, 102]. The hadronization mechanism employed is in many aspects very similar to the coalescence plus fragmentation approaches described in Sect. II.2.4. The coalescence for heavy mesons and baryons is evaluated in the Wigner formalism, see Eq. 6. The wave function is a Gaussian and the excited states are included assuming a suppression like in the statistical model as in the Catania model. Heavy quarks not recombining with light partons undergo fragmentation. Similarly to the SHM-TAMU approach, all the possible excited states are considered, including the unobserved baryon states not listed in the PDG, which are predicted to exist by the quark model and lattice QCD. A more “technical” difference is the mass of the light quarks which in EPOS are taken to be $m_{\mathrm{u}/d}=0.1\,\mathrm{G}eV$, while in the Catania, TAMU, and PHSD models values close to the constituent quark mass of about 0.33 GeV are used, reducing the violation of energy conservation in the coalescence process. Also EPOS4HQ has shown to generate for pp collisions ratios of $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ and $\Xi_{\mathrm{c}}^{0}/\mathrm{D^{0}}$ that are in fair agreement with the experimental data by ALICE and CMS, only if coalescence is included [101, 102]. Moreover, EPOS4HQ being a realistic fully dynamical approach showed that also the elliptic flow $v_{2}(p_{\mathrm{T}})$ of $\mathrm{D^{0}}$ in pp collisions agrees with the data and origins from the scattering of charm quarks in the medium generated in high-multiplicity events.
@@ III.1 Production in ee and ep collisions and spectroscopy
The precise theoretical knowledge of the total $\mathrm{c\overline{c}}$ cross section allows the direct measurement of the fragmentation fraction of charm quarks to a given charm-hadron species, $f(\mathrm{c\rightarrow H_{c}})$, without the need to measure the cross section of the other charm-hadron species. At LEP it was verified (see e.g. ALEPH analysis in Ref. [105]) that the total charm-production cross section, obtained as the sum of the measured $\mathrm{D^{0}}$, $\mathrm{D^{+}}$, $\mathrm{D_{s}^{+}}$, and $\Lambda_{\mathrm{c}}^{+}$ cross sections, and of an additional contribution for the unmeasured $\Xi_{\mathrm{c}}^{+,0}$ and $\Omega_{\mathrm{c}}^{0}$ cross sections, assumed to be around 2% of the measured one, is consistent with the Standard Model expectation. In Ref. [69] charm-hadron production measurements in $\mathrm{e^{+}e^{-}}$ collisions at B factories [110, 111, 112, 113, 114, 115, 116, 117] and in $\mathrm{Z^{0}}$-boson decays [106, 118, 105, 107, 119] were analysed together with data from photoproduction (PHP) and deep-inelastic scattering (DIS) in $\mathrm{e^{\pm}p}$ collisions at HERA [120, 121, 122, 123, 124], and with first results from pp collisions at the LHC. As shown in Fig. 8, no significant difference emerged among the charm FF measured in the various collision systems and collision energies. In $\mathrm{e^{\pm}p}$ and hadronic collisions charm-quark pairs are produced in hard-scattering processes from gluons, differently from $\mathrm{e^{+}e^{-}}$ collisions in which $\mathrm{g\rightarrow c\bar{c}}$ splittings occurs only in a fraction of events. At LEP, the fraction of $\mathrm{Z\rightarrow hadrons}$ events proceedings via $\mathrm{Z\rightarrow\mathrm{q\bar{q}g},g\rightarrow c\bar{c}}$ is about 3%, representing less than 20% of events with charm in Z decays [125, 126, 105, 127]. The consistency of the FF measured in $\mathrm{e^{+}e^{-}}$ and $\mathrm{e^{\pm}p}$ collisions was considered as a signal supporting the hypothesis of universality of the fragmentation fractions, which was expected to hold also in pp collisions at the LHC. First LHC data on D-meson production supported this hypothesis, which as will be discussed in next sections turned out not to be true.
@@
The probability that a given state is formed can depend on the hadron internal structure and state quantum numbers, for the determination of which spectroscopy measurements are essential. Charm and beauty baryons are also good probes to study diquark effective degrees of freedom. Theoretical calculations suggested that heavy-flavour baryons can be interpreted rather simply as a bound state of a diquark and a heavy quark [130, 133], a picture that requires experimental confirmation via a systematic comparison of measured and predicted spectra. Reviews on the topic, including discussions extended to exotic states, can be found in Ref. [130, 134, 135, 131, 136].
@@ III.2 Hadronic collisions: beam remnants and quantum-number conservation
Quantum-number conservation affects the relative abundances of the particle species produced at hadronization time leaving a trace that is more evident when phase-space for particle production is limited (e.g. hadronic collisions at low energies, particle decays) or in specific kinematic regions, e.g. close to beam rapidities. However, in high-energy hadronic collisions far from beam rapidities hadronization is not expected to be significantly influenced by the quark content of the beam particles. Of course, quantum-number conservation dictates a departure from the particle ratios observed in a “vacuum-like” system as $\mathrm{e^{+}e^{-}}$, and it spoils the effectiveness of factorization approach using $\mathrm{e^{+}e^{-}}$ FF in certain kinematic ranges. In many of these cases, the modelling of hadronization involves the recombination of valence quarks with other quarks produced in the event. Effects that can not be reconciled with a modelling of hadronization via universal fragmentation fractions/functions, were already observed at FNAL in $\pi$–A collisions. An example is the so-called beam-remnant drag [137], i.e. the tendency in a hadronic collision to produce at forward rapidity particles that share a valence quark (or diquark) with the remnant of the projectile moving in the same direction. Focusing only on the heavy-flavour sector, asymmetries in the production of $\mathrm{D^{-}}$/$\mathrm{D^{+}}$ [138, 139, 140, 141], $\mathrm{D_{s}^{-}}$/$\mathrm{D_{s}^{+}}$ [140, 142], $\Lambda_{\mathrm{c}}^{+}$/$\Lambda_{\mathrm{c}}^{-}$ [140, 143] and $\Lambda_{\mathrm{b}}^{0}$/$\bar{\Lambda}_{\mathrm{b}}^{0}$ [144] in various hadronic collisions have been observed. In Fig. 11 the asymmetry $A=\frac{\sigma_{\mathrm{D}^{-}}-\sigma_{\mathrm{D}^{+}}}{\sigma_{\mathrm{D}^{-}}+\sigma_{\mathrm{D}^{+}}}$ measured in $\pi^{\mathrm{-}}$–nucleus collisions (using platinum and diamond foils as target) by the E791 Collaboration is shown as a function of the Feynman $x_{\mathrm{F}}$ variable [141]. The positive asymmetry values emerging at large $x_{\mathrm{F}}$ indicate that the production of hadrons sharing valence quarks with beam hadron (the pion) is favoured.
@@ III.3 Proton–proton collisions
As several measurements show [151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161], $p_{\mathrm{T}}$-differential cross sections of open charm- and beauty-meson production in pp collisions are successfully described by quantum chromodynamics (QCD) calculations based on the factorization of the soft (non-perturbative) and hard (perturbative) processes and making use of FFs parameterized to reproduce measurements performed in $\mathrm{e^{+}e^{-}}$ or ep collisions. Implicitly, it is assumed in these calculations that the fragmentation and hadronization of charm and beauty quarks are universal processes independent of the collision system, or at least that they are approximately the same in $\mathrm{e^{+}e^{-}}$ and pp collisions. The charm-meson $p_{\mathrm{T}}$-integrated cross sections measured by several Collaborations in different colliding systems were used to compute the ratios of production yields among the different D-meson species ($\mathrm{D^{0}}$, $\mathrm{D^{+}}$, $\mathrm{D_{s}^{+}}$[162, 157, 163, 164, 165]. The prompt $\mathrm{D_{s}^{+}}$/($\mathrm{D^{0}}$ +$\mathrm{D^{+}}$) ratio represents the FF of charm quarks to charm-strange mesons $f_{\mathrm{s}}$ divided by the FF to non-strange charm mesons $f_{\mathrm{u}}+f_{\mathrm{d}},$ given that all $\mathrm{D^{*+}}$ and $\mathrm{D^{*0}}$ mesons decay to $\mathrm{D^{0}}$ and $\mathrm{D^{+}}$ mesons, and $\mathrm{D_{s}^{*+}}$ mesons decay to $\mathrm{D_{s}^{+}}$ mesons. The charm-quark fragmentation-fraction ratios $f_{\mathrm{s}}/(f_{\mathrm{u}}+f_{\mathrm{d}})$ from the ALICE [157, 166], H1 [124], ZEUS [120], and ATLAS [167] Collaborations are shown in the upper panel of Fig. 12. All the values are compatible within uncertainties and with the average of measurements at LEP [108]. The experimental data is also compared to the value obtained from a PYTHIA 8 simulation of pp collisions using the Monash tune [11], which is found to be compatible with the data within the uncertainties. The fragmentation fractions of beauty quarks to beauty-strange mesons divided by the one to non-strange beauty mesons computed by the ALICE [157], CDF [13], LHCb [168, 169], and ATLAS [170] Collaborations are also reported in the lower panel of Fig. 12. All the $f_{\mathrm{s}}/(f_{\mathrm{u}}+f_{\mathrm{d}})$ values measured in pp and $\mathrm{p\bar{p}}$ collisions are found to be compatible with the LEP average, computed by the HFLAV Collaboration [171], and with the value obtained from PYTHIA 8 simulation of pp collisions with the Monash tune. With a more recent measurement, the CMS Collaboration observed a mild $p_{\mathrm{T}}$ dependence ($\sim$ 10%) of $\mathrm{B}^{0}_{\mathrm{s}}/B^{0}$ ratio in the $p_{\mathrm{T}}$ interval 12–18 GeV$/c$ followed by a flat high-$p_{\mathrm{T}}$ trend, while the ratio $\mathrm{B}^{-}/B^{0}$ is found to be consistent with unity and independent of $p_{\mathrm{T}}$ and rapidity [172]. In general, the production cross sections and cross-section ratios of charm and beauty mesons do not show evidences of violation of the assumption of universality of the fragmentation and hadronization processes. However, recent measurements of charm and beauty baryon-to-meson production ratios and fragmentation fractions in pp collisions at LHC energies showed significant deviations from the values measured at $\mathrm{e^{+}e^{-}}$ and ep colliders, demonstrating that the assumption of universality of the hadronization process across collision systems has to be reconsidered.

### v3

@@ II.2.5 EPOS4(HQ)
Historically, the scattering modelling in EPOS was based on a zero-momentum limit of cut pomerons and hence did not aim at describing high-$p_{\perp}$ jet fragmentation (or other perturbative scattering processes). This has changed with EPOS4, which incorporates both factorization for hard processes and saturation for the low-$x$ scatterings. More specifically related to the purposes of this review, the main aspect of EPOS4HQ is the inclusion of the scattering dynamics of heavy quarks in the “corona” medium even in pp collisions and of hadronization via coalescence plus fragmentation [116, 117]. The hadronization mechanism employed is in many aspects very similar to the coalescence plus fragmentation approaches described in Sect. II.2.4. The coalescence for heavy mesons and baryons is evaluated in the Wigner formalism, see Eq. 6. The wave function is a Gaussian and the excited states are included assuming a suppression like in the statistical model as in the Catania model. Heavy quarks not recombining with light partons undergo fragmentation. Similarly to the SHM-TAMU approach, all the possible excited states are considered, including the unobserved baryon states not listed in the PDG, which are predicted to exist by the quark model and lattice QCD. A more “technical” difference is the mass of the light quarks which in EPOS are taken to be $m_{\mathrm{u}/d}=0.1\,\mathrm{G}eV$, while in the Catania, TAMU, and PHSD models values close to the constituent quark mass of about 0.33 GeV are used, which reduce the violation of energy conservation in the coalescence process. Furthermore, assuming a bulk system of quasi-particles with mass of about 0.1 GeV should lead to a entropy/pressure ratio close to $T_{c}$ quite larger than lattice QCD data, which are known to be well described by quasi-particle models with masses of about 0.3-0.4 GeV [118, 119]. Also in a coalescence approach a small mass with respect to the constituent one tends to favor baryon production especially at intermediate $p_{T}$ because it enhances more the probability to form baryons with respect to mesons due to the higher density and the larger boost effect of the relative momenta. A more thorough comparison analyzing how the results depend on the quark mass would be quite desirable.
EPOS4HQ has shown to generate for pp collisions ratios of $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ and $\Xi_{\mathrm{c}}^{0}/\mathrm{D^{0}}$ that are in fair agreement with the experimental data by ALICE and CMS, only if coalescence is included [116, 117]. Moreover, EPOS4HQ being a realistic fully dynamical approach showed that also the elliptic flow $v_{2}(p_{\mathrm{T}})$ of $\mathrm{D^{0}}$ in pp collisions agrees with the data and origins from the scattering of charm quarks in the medium generated in high-multiplicity events.
@@ III.1 Production in ee and ep collisions and spectroscopy
The precise theoretical knowledge of the total $\mathrm{c\overline{c}}$ cross section allows the direct measurement of the fragmentation fraction of charm quarks to a given charm-hadron species, $f(\mathrm{c\rightarrow H_{c}})$, without the need to measure the cross section of the other charm-hadron species. At LEP it was verified (see e.g. ALEPH analysis in Ref. [122]) that the total charm-production cross section, obtained as the sum of the measured $\mathrm{D^{0}}$, $\mathrm{D^{+}}$, $\mathrm{D_{s}^{+}}$, and $\Lambda_{\mathrm{c}}^{+}$ cross sections, and of an additional contribution for the unmeasured $\Xi_{\mathrm{c}}^{+,0}$ and $\Omega_{\mathrm{c}}^{0}$ cross sections, assumed to be around 2% of the measured one, is consistent with the Standard Model expectation. In Ref. [88] charm-hadron production measurements in $\mathrm{e^{+}e^{-}}$ collisions at B factories [127, 128, 129, 130, 131, 132, 133, 134] and in $\mathrm{Z^{0}}$-boson decays [123, 135, 122, 124, 136] were analysed together with data from photoproduction (PHP) and deep-inelastic scattering (DIS) in $\mathrm{e^{\pm}p}$ collisions at HERA [137, 138, 139, 140, 141], and with first results from pp collisions at the LHC. As shown in Fig. 8, no significant difference emerged among the charm FF measured in the various collision systems and collision energies. In $\mathrm{e^{\pm}p}$ and hadronic collisions charm-quark pairs are produced in hard-scattering processes from gluons, differently from $\mathrm{e^{+}e^{-}}$ collisions in which $\mathrm{g\rightarrow c\bar{c}}$ splittings occurs only in a fraction of events. At LEP, the fraction of $\mathrm{Z\rightarrow hadrons}$ events proceeding via $\mathrm{Z\rightarrow\mathrm{q\bar{q}g},g\rightarrow c\bar{c}}$ is about 3%, representing less than 20% of events with charm in Z decays [142, 143, 122, 144]. The consistency of the FF measured in $\mathrm{e^{+}e^{-}}$ and $\mathrm{e^{\pm}p}$ collisions was considered as a signal supporting the hypothesis of universality of the fragmentation fractions, which was expected to hold also in pp collisions at the LHC. First LHC data on D-meson production supported this hypothesis, which as will be discussed in next sections turned out not to be true.
@@
The probability that a given state is formed can depend on the hadron internal structure and state quantum numbers, for the determination of which spectroscopy measurements are essential. Charm and beauty baryons are also good probes to study diquark effective degrees of freedom. Theoretical calculations suggested that heavy-flavour baryons can be interpreted rather simply as a bound state of a diquark and a heavy quark [147, 150, 151], a picture that requires experimental confirmation via a systematic comparison of measured and predicted spectra. Reviews on the topic, including discussions extended to exotic states, can be found in Ref. [147, 152, 153, 148, 154].
@@ III.2 Hadronic collisions: beam remnants and quantum-number conservation
Quantum-number conservation affects the relative abundances of the particle species produced at hadronization time leaving a trace that is more evident when phase-space for particle production is limited (e.g. hadronic collisions at low energies, particle decays) or in specific kinematic regions, e.g. close to beam rapidities. However, in high-energy hadronic collisions far from beam rapidities hadronization is not expected to be significantly influenced by the quark content of the beam particles. Of course, quantum-number conservation dictates a departure from the particle ratios observed in a “vacuum-like” system as $\mathrm{e^{+}e^{-}}$, and it spoils the effectiveness of factorization approach using $\mathrm{e^{+}e^{-}}$ FF in certain kinematic ranges. In many of these cases, the modelling of hadronization involves the recombination of valence quarks with other quarks produced in the event. Effects that can not be reconciled with a modelling of hadronization via universal fragmentation fractions/functions, were already observed at CERN and FNAL in $\pi$–A collisions. An example is the so-called beam-remnant drag [155], i.e. the tendency in a hadronic collision to produce at forward rapidity particles that share a valence quark (or diquark) with the remnant of the projectile moving in the same direction. Focusing only on the heavy-flavour sector, asymmetries in the production of $\mathrm{D^{-}}$/$\mathrm{D^{+}}$ [156, 157, 158, 159], $\mathrm{D_{s}^{-}}$/$\mathrm{D_{s}^{+}}$ [158, 160], $\Lambda_{\mathrm{c}}^{+}$/$\Lambda_{\mathrm{c}}^{-}$ [158, 161] and $\Lambda_{\mathrm{b}}^{0}$/$\bar{\Lambda}_{\mathrm{b}}^{0}$ [162] in various hadronic collisions have been observed. In Fig. 11 the asymmetry $A=\frac{\sigma_{\mathrm{D}^{-}}-\sigma_{\mathrm{D}^{+}}}{\sigma_{\mathrm{D}^{-}}+\sigma_{\mathrm{D}^{+}}}$ measured in $\pi^{\mathrm{-}}$–nucleus collisions (using platinum and diamond foils as target) by the E791 Collaboration is shown as a function of the Feynman $x_{\mathrm{F}}$ variable [159]. The positive asymmetry values emerging at large $x_{\mathrm{F}}$ indicate that the production of hadrons sharing valence quarks with beam hadron (the pion) is favoured.
@@ III.3 Proton–proton collisions
As several measurements show [169, 170, 171, 172, 173, 174, 175, 19, 176, 177, 178], $p_{\mathrm{T}}$-differential cross sections of open charm- and beauty-meson production in pp collisions are successfully described by quantum chromodynamics (QCD) calculations based on the factorization of the soft (non-perturbative) and hard (perturbative) processes and making use of FFs parameterized to reproduce measurements performed in $\mathrm{e^{+}e^{-}}$ or ep collisions. Implicitly, it is assumed in these calculations that the fragmentation and hadronization of charm and beauty quarks are universal processes independent of the collision system, or at least that they are approximately the same in $\mathrm{e^{+}e^{-}}$ and pp collisions. The charm-meson $p_{\mathrm{T}}$-integrated cross sections measured by several Collaborations in different colliding systems were used to compute the ratios of production yields among the different D-meson species ($\mathrm{D^{0}}$, $\mathrm{D^{+}}$, $\mathrm{D_{s}^{+}}$[179, 175, 180, 181, 182]. The prompt $\mathrm{D_{s}^{+}}$/($\mathrm{D^{0}}$ +$\mathrm{D^{+}}$) ratio represents the FF of charm quarks to charm-strange mesons $f_{\mathrm{s}}$ divided by the FF to non-strange charm mesons $f_{\mathrm{u}}+f_{\mathrm{d}},$ given that all $\mathrm{D^{*+}}$ and $\mathrm{D^{*0}}$ mesons decay to $\mathrm{D^{0}}$ and $\mathrm{D^{+}}$ mesons, and $\mathrm{D_{s}^{*+}}$ mesons decay to $\mathrm{D_{s}^{+}}$ mesons. The charm-quark fragmentation-fraction ratios $f_{\mathrm{s}}/(f_{\mathrm{u}}+f_{\mathrm{d}})$ from the ALICE [175, 183], H1 [141], ZEUS [137], and ATLAS [184] Collaborations are shown in the upper panel of Fig. 12. All the values are compatible within uncertainties and with the average of measurements at LEP [125]. The experimental data is also compared to the value obtained from a PYTHIA 8 simulation of pp collisions using the Monash tune [11], which is found to be compatible with the data within the uncertainties. The fragmentation fractions of beauty quarks to beauty-strange mesons divided by the one to non-strange beauty mesons computed by the ALICE [175], CDF [29], LHCb [185, 13], and ATLAS [186] Collaborations are also reported in the lower panel of Fig. 12. All the $f_{\mathrm{s}}/(f_{\mathrm{u}}+f_{\mathrm{d}})$ values measured in pp and $\mathrm{p\bar{p}}$ collisions are found to be compatible with the LEP average, computed by the HFLAV Collaboration [187], and with the value obtained from PYTHIA 8 simulation of pp collisions with the Monash tune. The LHCb and CMS Collaborations observed a mild $p_{\mathrm{T}}$ dependence ($\sim$ 10%) of $f_{\mathrm{s}}/(f_{\mathrm{u}}$ ratio in the low/intermetdiate $p_{\mathrm{T}}$ region followed by a flat high-$p_{\mathrm{T}}$ trend, while the ratio $f_{\mathrm{d}}/(f_{\mathrm{u}}$ is found to be consistent with unity and independent of $p_{\mathrm{T}}$ and rapidity [188, 189]. In general, the production cross sections and cross-section ratios of charm and beauty mesons do not show evidences of violation of the assumption of universality of the fragmentation and hadronization processes. However, recent measurements of charm and beauty baryon-to-meson production ratios and fragmentation fractions in pp collisions at LHC energies showed significant deviations from the values measured at $\mathrm{e^{+}e^{-}}$ and ep colliders, demonstrating that the assumption of universality of the hadronization process across collision systems has to be reconsidered.